On your cousin’s wedding day you leave for the church 30.0 min before the ceremony is to begins, which should leave plenty of time since the church is only 40.0 km away. On the way, however, you have to make an unanticipated stop because of road construction work. As a result, your average speed for the first 10 min is only 15.0 km/h.


What average speed do you need to have on the remainder of the trip to get you to the church on time?

1
Expert's answer
2022-03-29T12:27:26-0400

"s_1=v_1t_1=2.5~km,"

"s_2=s-s_1=27.5~km,"

"t_2=t-t_1=30~min,"

"v_2=\frac{s_2}{t_2}=55~\frac{km}h."